# From Internal Conflict to Contextual Adaptation of Language Models

### v1

@@ Abstract
Knowledge-intensive language understanding tasks require Language Models (LMs) to integrate relevant context, mitigating their inherent weaknesses, such as incomplete or outdated knowledge. Nevertheless, studies indicate that LMs often ignore the provided context as it can conflict with the pre-existing LM’s memory learned during pre-training. Moreover, conflicting knowledge can already be present in the LM’s parameters, termed intra-memory conflict. Existing works have studied the two types of knowledge conflicts only in isolation. We conjecture that the (degree of) intra-memory conflicts can in turn affect LM’s handling of context-memory conflicts. To study this, we introduce the DynamicQA dataset, which includes facts with a temporal dynamic nature where a fact can change with a varying time frequency and disputable dynamic facts, which can change depending on the viewpoint. DynamicQA is the first to include real-world knowledge conflicts and provide context to study the link between the different types of knowledge conflicts. With the proposed dataset, we assess the use of uncertainty for measuring the intra-memory conflict and introduce a novel Coherent Persuasion (CP) score to evaluate the context’s ability to sway LM’s semantic output. Our extensive experiments reveal that static facts, which are unlikely to change, are more easily updated with additional context, relative to temporal and disputable facts.
@@ 1 Introduction
Language models (LMs) have been useful in a variety of downstream applications from summarization to fact-checking, often relying on the factual knowledge memorized during pre-training and stored in their parameters, known as parametric memory Yu et al. (2024). However, this internal memory is not infallible; it may contain misinformation, biases or simply outdated data, causing LMs to produce factually incorrect output, occasionally termed ‘hallucinations’ Huang et al. (2023). One way to mitigate this issue is the use of retrievers Guu et al. (2020) to provide additional context to a language model (non-parametric knowledge). However, this context may conflict with the model’s parametric knowledge, resulting in a knowledge conflict.
Handling these knowledge conflicts has been a productive area for research. Initial research looked at behavioural patterns within LMs, where they found a tendency of models to over-rely on parametric knowledge in the face of these knowledge conflicts Longpre et al. (2021); Chen et al. (2022), thus ignoring the retrieved context. Conflicts triggered by the contradiction of a context and a model’s parametric knowledge is often termed context-memory conflicts. Context that conflicts with a model’s parametric knowledge is often termed counter-memory Xie et al. (2024).
Figure 1: We present examples of our dataset here, consisting of static, temporal, disputable facts. It shows how the answers to the question can be different depending on the types of facts.
Many of these studies overlook that knowledge conflicts can also happen within the parametric knowledge, for example, due to conflicting representations of the fact in the training data. These type of knowledge conflicts, called intra-memory conflicts, may contribute to LMs’ uncertainty and instability Marjanović et al. (2024). Furthermore, we argue that the existence of intra-memory conflicts should impact LM’s behaviour in context-memory conflicts. Ideally, the degree of conflict in parametric memory should determine how much the model relies on non-parametric memory. While intra-memory conflicts remain understudied, elucidation of their interaction with context-memory conflicts impacts work within retrieval augmentation and knowledge editing.
In this paper, we study knowledge conflicts in relation to fact dynamicity to effectively simulate realistic knowledge conflicts. To facilitate the approach, we introduce a new dataset for knowledge conflict, DynamicQA, which contains facts with presumed entropy in our pretraining dataset alongside static facts, as a control. DynamicQA consists of three types of facts: static, temporal, and disputable. Temporal facts are facts that are changed by time and disputable facts depend on one’s viewpoint (Figure 1). Furthermore, we introduce proxy scores to approximate the dynamicity of these facts. We then assess the use of semantic entropy Kuhn et al. (2023) as an indicator of intra-memory conflict, and introduce a novel Coherent Persuasion (CP) score to measure new context’s ability to change a LM’s semantic output.
@@
With DynamicQA, we find that manipulated contexts of static facts and temporal facts with low temporality show the greatest persuasiveness, despite their limited historical change and, presumably, future likelihood to change. Furthermore, we find that semantic uncertainty is a poor indicator of intra-memory conflict and it alone does not reflect an instance’s likelihood for persuasion given additional context, suggesting that retrieval-augmented generation language models struggle in low-certainty domains. These results underline the need for new measures of intra-memory conflict in the absence of our provided proxy scores and other indicators of adaptive-retrieval efficiency Ni et al. (2024).
@@ 2 Related Work
Xu et al. (2024b) provide a survey for knowledge conflicts in LMs; we focus specifically on intra-memory conflicts and context-memory conflicts. Longpre et al. (2021) first showed that LMs ignore context that contradicts their parametric memory, generating facts not stated in the input, termed as ‘hallucinations’. Further work either disentangles the two forms of knowledge conflicts Neeman et al. (2023), or finds other methods to resolve LM hallucination Song et al. (2024). Other studies investigate the effect of a fact’s frequency in the training data Yu et al. (2023) or the quality of the presented counter-memory Xie et al. (2024); Wan et al. (2024) on LM hallucination. Furthermore, some works find that LMs can be swayed by convincing misinformation at inference-time Pan et al. (2023); Xu et al. (2024a); Wan et al. (2024). The existing findings suggest some pattern to a model’s proclivity to ignore context. However, in contrast to these existing works, our work considers realistic, natural examples of knowledge conflict as they occur in the real world and how different types of knowledge conflicts are connected.
@@
There have been several approaches to quantify or approximate the degree of a knowledge conflict; however, these works only focus on context-memory conflicts. Pezeshkpour (2023) focus on the entropy of the LM’s probability distribution for an answer. The difference in entropy depending on the presence or absence of additional context indicates the LM’s prior knowledge on a certain fact. On the other hand, Du et al. (2024) propose a susceptibility and persuasion score to investigate the LM when the given context contradicts the LM’s parametric knowledge, based on the first token probabilities. The susceptibility score of an entity shows how easy it is to shift an LM’s probability distribution of an answer regarding an entity. The persuasion score of a context represents how effective the context is at changing an LM’s probability distribution of an answer. Given known issues in the reliance on first-token probabilities Wang et al. (2024), we reformulate this score with consideration of semantic consistency in §3.3. Furthermore, given the lack of measures for intra-memory conflict, we assess semantic entropy Kuhn et al. (2023) as an indicator of intra-memory conflict Gao et al. (2024).
@@
We identify three causes of model factual error: fact popularity, temporality, and disputability. Two of these causes have been previously investigated in other general studies. Fact popularity approximates the prevalence of a fact in training data and has been shown to affect model performance Mallen et al. (2023) and a model’s likelihood to be swayed by context (‘susceptibility score’, Du et al. (2024)). In contrast, temporality and disputability reflect a fact’s entropy within the training and test data. For example, it is well-established that the passing of time can lead to outdated models, stressing the need for evaluation of models for a specific time-frame Margatina et al. (2023); Fierro et al. (2024) and updating of specific facts in LMs Zhang and Choi (2023); Jang et al. (2022). Similarly, Xu et al. (2024b) argue that context-memory knowledge conflicts arise due to either temporal misalignment or misinformation pollution of the training dataset. In a contemporary study, Fierro et al. (2024) show that facts with dynamic (or in their terms ‘mutable’) relations show different patterns in model confidence, knowledge representation, and contextual alignment. All of these studies look at only one source of factual error at a time and do not assess in the context of knowledge conflicts. In our work, we introduce disputability as a contributor to factual error, study multiple sources of factual error concurrently (temporality, disputability, and popularity), assess their interaction, and provide real-world proxy measures for all three sources.
@@ 3 Measuring Knowledge Conflicts
We describe our two proposed methods to measure knowledge conflicts: semantic uncertainty Kuhn et al. (2023) and our novel Coherent Persuasion score. Both measures look primarily at semantic changes in LM output, thereby limiting confounds due to syntactic inconsistencies and other issues arising from over-reliance on first-token probabilities. We first describe semantic entropy as a reflection of the intra-memory conflict, and how we determine semantic entailment (§3.2) and then present our Coherent Persuasion (CP) score as a measure of context-memory conflict (§3.3).
@@ 3.1 Preliminaries
A dataset $X=[(c_{1},q_{1},y_{1}),\dots,(c_{N},q_{N},y_{N})]$ consists of $N$ tuple instances containing: an answer-containing context $c$, a question $q$, and a ground-truth answer $y$. For the $i$-th instance, an LM $f$ outputs an answer $a_{i}=[a_{i}^{1},\dots,a_{i}^{h},\dots,a_{i}^{H}]$, consisting of $H$ number of tokens given an input $x_{i}$. An input $x_{i}$ here contains a universal prompt $P$ describing the QA task with either $q_{i}$ only ($x_{i}=[P;q_{i}]$) or with $q_{i}$ and the context $c_{i}$ ($x_{i}=[P;c_{i};q_{i}]$).
3.2 Semantic Uncertainty Score
Uncertainty can arise in question-answering tasks as several possible related facts may exist within an LM’s parametric knowledge, especially in cases of high intra-memory conflict Xu et al. (2024b). Given its use of high-temperature sampling to elicit diverse answers contained within parametric memory, we propose the use of semantic uncertainty as an estimator of intra-memory conflict. We then validate this measure using its relation to a fact’s dynamicity (defined in §4).
To do so, we use Kuhn et al. (2023)’s approach in our simulated knowledge conflict setting, which we then adapt for our Persuasion Score in §3.3. Given the input $x_{i}$, which contains a universal prompt and a question $q_{i}$, we first generate $K$ number of answers $A=[a_{i,0},a_{i,1},\dots,a_{i,K}]$. Next, we group the generated answers according to their semantic similarity. The semantic similarity between two sampled answers is calculated using a DeBERTA Natural Language Inference (NLI) model. The details about answer generation and semantic grouping are in Appendix A.3. The grouping according to the semantic similarity results in the $V$ number of groups $G=[g_{1},g_{2},\dots,g_{v},\dots,g_{V}]$, where $V$ can range from 1 to $K$.
The semantic uncertainty is then estimated by the entropy between the semantic sets. First, we obtain $p(g_{v}|q_{i})$, the probability of the model generating the answers in $g_{v}$ as in Equation 1.
${}\begin{split}p(g_{v}|x_{i})=\sum_{a_{i,k}\in g_{v}}p(a_{i,k}|x_{i})\\ =\sum_{a_{i,k}\in g_{v}}\prod_{h}p(a_{i}^{h}|a_{i}^{<h},x_{i})\end{split}$ (1)
With the probabilities of groups, we approximate $SE(x_{i})$, the expectation of the semantic entropy of $x_{i}$, using Monte Carlo integration over the groups:

### v2

# DynamicQA: Tracing Internal Knowledge Conflicts in Language Models

@@ Abstract
Knowledge-intensive language understanding tasks require Language Models (LMs) to integrate relevant context, mitigating their inherent weaknesses, such as incomplete or outdated knowledge. However, conflicting knowledge can be present in the LM’s parameters, termed intra-memory conflict, which can affect a model’s propensity to accept contextual knowledge. To study the effect of intra-memory conflict on an LM’s ability to accept relevant context, we utilize two knowledge conflict measures and a novel dataset containing inherently conflicting data, DynamicQA. This dataset includes facts with a temporal dynamic nature where facts can change over time and disputable dynamic facts, which can change depending on the viewpoint. DynamicQA is the first to include real-world knowledge conflicts and provide context to study the link between the different types of knowledge conflicts. We also evaluate several measures on their ability to reflect the presence of intra-memory conflict: semantic entropy and a novel coherent persuasion score. With our extensive experiments, we verify that LMs exhibit a greater degree of intra-memory conflict with dynamic facts compared to facts that have a single truth value. Furthermore, we reveal that facts with intra-memory conflict are harder to update with context, suggesting that retrieval-augmented generation will struggle with the most commonly adapted facts.
@@ 1 Introduction
Language models (LMs) have been useful in a variety of downstream applications from summarization to fact-checking, often relying on the factual knowledge memorized during pre-training and stored in their parameters, known as parametric knowledge Yu et al. (2024). However, this internal memory is not infallible; it may contain misinformation, biases or simply outdated data, causing LMs to produce factually incorrect output, occasionally termed ‘hallucinations’Huang et al. (2023). Furthermore, conflicting representations of a fact may exist within training data, given conflicting viewpoints (‘disputable facts’) or temporal changes (‘temporal facts’) (See Figure 1). We refer to this superposition of fact representation as the fact’s dynamicity. In contrast, a static fact has only one possible representation. This can lead to intra-memory conflicts, which can contribute to LM’s uncertainty and instability in factual recall. This intra-memory conflict can be likened to a semantic expression of aleatoric uncertainty, noise inherent in training data Kendall and Gal (2017).
One way to guide LMs to reliable answers is via Retrieval-Augmented Generation (RAG) Guu et al. (2020), which provides additional contextual knowledge, known as non-parametric knowledge, to a language model. However, this context may conflict with the model’s parametric knowledge, resulting in a context-memory conflict. Initial research has found a tendency of models to over-rely on parametric knowledge in the face of these knowledge conflicts Longpre et al. (2021); Chen et al. (2022), thus ignoring the retrieved context. While fact popularity has been known to reduce context utilization Mallen et al. (2023), little is known of the impact of internal memory conflicts. We explore this issue for the first time by looking at the effect of fact dynamicity on context utilization.
Figure 1: We present examples of our dataset here, consisting of static, temporal, and disputable facts. We show how model output distribution can vary due to the popularity (a,b) and dynamicity (c) of facts. Fact dynamicity (c) causes intra-memory conflicts between the different fact representations seen during pretraining. In the bottom row, we show the change in output probability (purple area) that the context must enact on the initial output distribution (dotted line) to force a new output distribution (purple line).
In this paper, we investigate the effect of intra-memory conflicts on context adaptation using a novel dataset DynamicQA. This dataset contains inherently conflicting temporal and disputable facts, paired with realistic retrieved contexts in which they occur. Furthermore, we manipulate these contexts to include similar, competing objects. Using DynamicQA, we study the interaction of intra-memory and context-memory conflicts by adapting two existing conflict measures. First, we approximate an LM’s fact dynamicity ground truth via the number of recorded representations for a fact in Wikidata, and sample the LM’s expressed intra-memory conflict using semantic entropy, which captures the semantic variation present in parametric memory Kuhn et al. (2023). Furthermore, we introduce a coherent persuasion score, based on Du et al. (2024) to approximate an LM’s semantic shift in output distribution given competing context.
@@
With our framework, we show the effect of intra-memory conflict on dynamic fact question answering. We find that manipulated contexts of static facts and facts with low dynamicity show the greatest persuasiveness for LMs, despite their limited variance in the training dataset and, presumably, low likelihood to change. Furthermore, we find that semantic entropy alone does not reflect an instance’s likelihood for persuasion given additional context, suggesting that retrieval-augmented generation language models struggle in low-certainty domains. These results underline the need for new measures of intra-memory conflict and the need for other indicators of context use in RAGs Ni et al. (2024).
@@ 2 Related Work
Xu et al. (2024b) provide a survey for knowledge conflicts in LMs; we focus specifically on intra-memory conflicts and their interplay with context-memory conflicts. Longpre et al. (2021) first showed that LMs ignore context that contradicts their parametric memory, generating facts not stated in the input, termed ‘hallucinations’. Further work either disentangles the two forms of knowledge Neeman et al. (2023) or finds other methods to resolve hallucination Song et al. (2024). Other studies investigate the effect of a fact’s frequency in the training data Yu et al. (2023) or the quality of the presented counter-memory Xie et al. (2024); Wan et al. (2024) on LM hallucination. Furthermore, some works find that LMs are swayed by convincing misinformation at inference-time Pan et al. (2023); Xu et al. (2024a); Wan et al. (2024). These findings suggest some pattern to a model’s proclivity to ignore context. However, in contrast to these existing works, our work considers realistic, natural examples of knowledge conflict as they occur in the real world and how different types of knowledge conflicts are connected.
@@
There have been several approaches to quantify or approximate the degree of a knowledge conflict; however, these works only focus on context-memory conflicts. Pezeshkpour (2023) focus on the entropy of the LM’s probability distribution for an answer; the difference in entropy depending on the presence or absence of additional context indicates the LM’s prior knowledge on a certain fact. On the other hand, Du et al. (2024) propose a susceptibility and a persuasion score to investigate the LM when the given context contradicts the LM’s parametric knowledge, based on the first token probabilities. The susceptibility score of an entity shows how easy it is to shift an LM’s probability distribution of an answer regarding an entity. The persuasion score of a context represents how effective the context is at changing an LM’s probability distribution of an answer. Given known issues in the reliance on first-token probabilities Wang et al. (2024), we reformulate this score with consideration of semantic consistency in §4.3. Furthermore, given the lack of measures for intra-memory conflict, we assess semantic entropy Kuhn et al. (2023) as an indicator of intra-memory conflict Gao et al. (2024).
@@
We identify three causes of model factual error: fact popularity, temporality, and disputability. Fact popularity approximates the prevalence of a fact in training data and has been shown to affect model performance Mallen et al. (2023) and a model’s likelihood to be swayed by context (‘susceptibility score’, Du et al. (2024)). In contrast, temporality and disputability reflect a fact’s entropy within the training and test data. For example, it is well-established that the passing of time can lead to outdated models, stressing the need for evaluation of models for a specific time-frame Margatina et al. (2023); Fierro et al. (2024) and updating of specific temporal facts in LMs Zhang and Choi (2023); Jang et al. (2022). Similarly, Xu et al. (2024b) argue that context-memory knowledge conflicts arise due to either temporal misalignment or misinformation pollution of the training dataset. In a contemporary study, Fierro et al. (2024) show that facts with dynamic (or in their terms ‘mutable’) relations show different patterns in model confidence, knowledge representation, and contextual alignment. All of these studies look at only one source of factual error at a time and do not assess in the context of knowledge conflicts. In our work, we introduce disputability as a contributor to factual error, study multiple sources of factual error concurrently (temporality, disputability, and popularity), assess their interaction, and provide real-world proxy measures for all three sources.
3 DynamicQA
To evaluate the intra-memory conflicts, we present a dataset of 11,378 question-answer pairs featuring facts with varying levels of dynamicity (see Figure 1). The dynamicity of a fact is difficult to determine via the pretraining dataset, so we use easily measured proxy scores as a reflection of the property. The question and answer pairs, alongside their proxy scores, are sourced from Wikidata and Wikipedia and have not previously been employed in similar datasets nor for approximating the degree of knowledge conflicts in general. We approximate temporality via the number of edits (§3.1) and disputability via the number of reversions (§3.2). For each question and answer pair, we obtain answer-specific context snippets from Wikipedia, alongside their popularity scores, estimated from Wikipedia page views. For static and temporal facts, we also create ‘counter-memory’ from these snippets using similar object replacements, to simulate realistic knowledge conflict scenarios.
3.1 Temporality
We estimate the temporality of a triplet via the number of edits on Wikidata for the given subject and relation as a proxy. We initialize our dataset from PopularQA Mallen et al. (2023), which is a collection of 14k question-answer pairs sampled from Wikipedia, with a popularity score determined by the monthly Wikipedia page views of the subject and object of the triplets. This popularity score is often used as a proxy for the triplet’s prevalence in the unsearchable pre-training corpora of models Mallen et al. (2023); Fierro et al. (2024). Given the sampling method, this dataset has a long-tailed distribution, meaning most triplets have very low popularity and low temporality. We identify relevant snippets from the triplet’s subject current Wikipedia page that also mention the triplet’s object and use them as context to provide to the LM. We discard 2k pairs where we cannot find the intended object and relation mentioned on the subject’s Wikipedia page. To construct counter-memory contexts, we replace the object in context with a replacement token. We identify relevant replacement entities using the most similar entity for the original object as identified using the Wembedder tool Årup Nielsen (2017). All facts with more than 1 edit are labelled as ‘temporal’ facts, though they vary in their degree of temporality (with a maximum score of 23 edits). This leaves us with 2495 questions (Table 1). We randomly subsample 2500 of the remaining ‘static’ questions (triplets with 0 edits on Wikidata).
3.2 Disputability
Disputable facts are facts that differ depending on the viewpoint. To construct the question and answer pairs, we utilize the collection of Wikipedia articles that are regarded as controversial. [^1] https://en.wikipedia.org/wiki/Wikipedia:List_of_controversial_issues Controversy, by definition, is inherently disputable since it arises from the existence of multiple viewpoints, each supported by sufficient evidence Rad and Barbosa (2012); Vuong et al. (2008). We are the first to convert the controversies that appear in the Wikipedia edit history into a QA dataset for LMs.
To identify disputable facts from the controversial articles, we look for reversions in the edit logs Sumi et al. (2011). Among the edit logs $[e_{0},\dots,e_{l-1},e_{l},e_{l+1},\dots,e_{j}]$ on an article, two consecutive logs, $e_{l}$ and $e_{l+1}$ are selected as a pair of reverted edit logs if the texts of $e_{l-1}$ and $e_{l+1}$ are identical. Then, by measuring the edit distance between $e_{l}$ and $e_{l+1}$, we select a pair of words from $e_{l}$ and $e_{l+1}$ if replacing the word in $e_{l}$ with the word in $e_{l+1}$ yields the same text as $e_{l+1}$.
Since a reverted edit does not necessarily imply the disputability of a fact, we apply additional rules to filter out cases of vandalism, paraphrasing, or synonyms. For vandalism, we remove the pair if one of the users involved in that edit did not disclose their identity, for example, anonymous users or if the user ID is an IP address. For the pairs that are synonyms or paraphrasing, we feed them to a semantic similarity model (§A.3) and remove the pairs whose similarity scores are bigger than 0.98. The selected pairs of reverted edits and the text snippets they are located in serve correspondingly as answers and contexts in our dataset.
Questions are generated using an LM [^2] https://huggingface.co/meta-llama/Meta-Llama-3-8B-InstructA.2) by providing the context $e_{l}$ and the corresponding ground-truth answer within the prompt (see Table 5 in the Appendix). After obtaining the question and answer pair with the context, three annotators manually annotated the dataset to ensure its quality. Two annotators annotated each data point, and conflicts were resolved by the third annotator. We obtain a Krippendorf’s alpha of 0.44 and provide further details in Appendix A.3. For the real-world proxy score of disputability, we count the number of reversions on $(e_{l},e_{l+1})$. As a result, we obtain 694 questions with two possible disputable answers and their corresponding real-world proxy scores (Table 1).
@@ 4 Measuring Knowledge Conflicts
In the pretraining stage, LMs learn a conditional distribution to predict the next tokens given a preceding context; in this process, they also obtain world knowledge into their parametric memory Chang et al. (2024). For example, in Figure 1, given a question (e.g., ‘Who is the father of Queen Elizabeth II?’), the conditional probability of the correct answer ‘George IV’ is expected to be greater than similar objects, owing to the frequent co-occurrence of the subject, object and relation in the training corpus. One important element impacting fact acquisition and recall is the frequency of the fact in the training data. It can be assumed that high frequency (often approximated via fact popularity Mallen et al. (2023)) leads to a greater strength of association between the subject and the object, resulting in greater probability being attributed to the relevant tokens (Figure 1b). However, as information evolves over time, multiple different representations of a factual triplet may appear in a training corpus; meaning, given a subject and relation, there can exist a superposition of object representations Shanahan et al. (2023) depending on the temporal and situational context. Rather than a singular object being assigned a high probability, for such facts, we expect several competing answers in the output distribution (Figure 1c).
To evaluate this hypothesis, we expect to see two properties: (1) higher entropy in the output distribution of dynamic facts and (2) greater change in the output distribution to update dynamic facts. To verify (1), we look at the semantic entropy of output distributions (§4.2). To investigate (2), we analyse the amount of effort required to shift the output distribution using both loss and our novel Coherent Persuasion score (§4.3). Both measures look at semantic changes in LM output, thereby limiting confounds due to syntactic inconsistencies and other issues arising from over-reliance on first-token probabilities Wang et al. (2024).
@@ 4.1 Preliminaries
A dataset $D=[(c_{1},q_{1},y_{1}),\dots,(c_{N},q_{N},y_{N})]$ consists of $N$ tuple instances containing: an answer-containing context $c$, a question $q$, and a ground-truth answer $a$. For the $i$-th instance, an LM $f$ outputs an answer $y_{i}=[y_{i}^{1},\dots,y_{i}^{h},\dots,y_{i}^{H}]$, consisting of $H$ tokens given an input $x_{i}$. An input $x_{i}$ here contains a universal prompt $P$ describing the QA task with either $q_{i}$ only ($x_{i;q}=[P;q_{i}]$) or with $q_{i}$ and the context $c_{i}$ ($x_{i;q,c}=[P;q_{i},c_{i}]$). Given a contextless input ($x_{i;q}$), a model outputs $y_{i;q}$ (i.e. the parametric knowledge). Given context ($x_{i;q,c}$), we denote the model output as $y_{i;q,c}$.
4.2 Semantic Entropy
Entropy allows us to measure the amount of information within a distribution and is a common measure for model uncertainty. Both lack of information (Figure 1a) and competing information (Figure 1c) can contribute to uncertainty and lead to more uniform output distributions. We propose semantic entropy as an estimator of intra-memory conflict, given its use of high-temperature sampling to elicit diverse answers contained within parametric memory. We validate this measure through the lens of a fact’s dynamicity (defined in §3).
To do so, we use Kuhn et al. (2023)’s approach in our simulated knowledge conflict setting. Given any input $x_{i}$, which contains a universal prompt and a question $q_{i}$ (and optional context $c_{i}$), we generate $K$ model outputs $Y=[y_{i,1},\dots,y_{i,K}]$. Next, we group the generated answers according to their semantic similarity. The semantic similarity between two sampled answers is calculated using a DeBERTA Natural Language Inference (NLI) model. [^3] https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 Details about answer generation and semantic grouping can be found in Appendix A.3. The grouping according to the semantic similarity results in the $V$ groups $G=[g_{1},g_{2},\dots,g_{v},\dots,g_{V}]$, where $1\leq V\leq K$.
The semantic entropy is then estimated by the entropy between the semantic sets. First, we obtain $p(g_{v}|x_{i})$, the conditional token probabilities output by the model generating the answers in $g_{v}$ given the input $x_{i}$ via Equation 1.
${}\begin{split}p(g_{v}|x_{i})=\sum_{y_{i,k}\in g_{v}}p(y_{i,k}|x_{i})\\ =\sum_{y_{i,k}\in g_{v}}\prod_{h}p(y_{i}^{h}|y_{i}^{<h},x_{i})\end{split}$ (1)
In this case, $h$ refers to the intermediate tokens in the entire output length. With the probabilities of the separate groups, we approximate $SE(x_{i})$, the overall expectation of the semantic entropy of $x_{i}$, using Monte Carlo integration over the groups: